# What Is a Good Caption? A Comprehensive Visual Caption Benchmark for Evaluating Both Correctness and Thoroughness

### v2

@@ Methods
F1-score LLaVA-OV-72B 78.7 36.5 67.1 62.2 89.7 51.7 76.1 83.7 16.8 41.6 41.1 14.4 43.4 54.1
Qwen2VL-72B 81.0 36.9 69.8 64.3 89.9 52.3 87.0 82.6 10.7 29.7 43.1 29.0 46.4 55.6
InternVL2.5-78B 79.0 40.2 70.5 63.8 89.6 48.4 85.3 82.8 20.5 32.7 34.1 14.7 38.8 53.9
Qwen2.5VL-72B 81.8 40.3 70.0 67.8 91.1 54.2 91.2 84.7 34.5 35.8 44.6 35.0 49.2 60.0
GPT-4o (0806) 85.5 41.5 74.6 68.9 89.9 67.0 87.9 90.0 41.7 43.8 53.1 26.6 56.7 63.6
Gemini-1.5-pro 88.0 51.5 76.5 72.5 89.2 56.4 89.9 91.4 43.5 60.4 54.0 34.6 59.9 66.8
Gemini-2.0-flash 84.1 44.1 72.7 65.5 89.6 57.4 92.0 86.8 51.9 44.0 48.5 35.3 54.9 63.6
3.2 Multiple Dimension Evaluation
Caption evaluation. As we collect and annotate the data of each dimension separately, we also evaluate each dimension independently. Different from matching the similarity between the caption and ground-truth sentences, our annotation drops the caption sentence, and we use GPT-4 Turbo [1] to take interpretable scores for each dimension. We use a similar prompt template for dimensions with specific categories (i.e., style, camera angle, and camera movement), and use another similar prompt template for other open-set dimensions. See Sec. B.2 for the details of the prompts. Therefore, we can judge the caption into the following three situations:
  • Miss, which means the caption does not mention the corresponding content about the dimension.
  • Positive, which means the caption mentions the corresponding content about the dimension, and describes it correctly compared with the annotations.
  • Negative, which means the caption mentions the corresponding content about the dimension, but gives a wrong description compared with the annotations.
As all data can be judged into these three situations, we can then calculate two metrics: 1) Precision, which represents the accuracy on all samples that the model has mentioned, and thus only considers the precision. 2) Recall, which represents the accuracy on all annotated samples, no matter whether the dimension is described or missed in the caption, and thus considers both the correctness and thoroughness. Given the set of all samples as $S(\text{All})$, positive samples as $S(\text{Pos})$, negative samples as $S(\text{Neg})$, missed samples as $S(\text{Miss})$, the metrics can be calculated by:
Precision $=\frac{|S(\text{Pos})|}{|S(\text{All})-S(\text{Miss})|},$ (1)
Recall $=\frac{|S(\text{Pos})|}{|S(\text{All})|}.$ (2)
We report their harmonic mean (i.e., F1-score) as our main metric. Apart from them, we also introduce a new metric, Hit rate, which represents the referring ratio about the dimension in visual caption and can be calculated as:
Hit $\text{Hit rate}=\frac{|S(\text{All})-S(\text{Miss})|}{|S(\text{All})|}.$ (3)
This metric only considers the pure thoroughness of the caption in each dimension, without thinking about the accuracy.
@@ 3.2 Multiple Dimension Evaluation
Question-answer pairs evaluation. As we annotate each descriptive element for each dimension rather than caption sentence, we can also convert our annotations to question-answer (QA) pair format to evaluate the MLLMs’ general ability out of the horizon of caption only. See Sec. B.4 for examples of our CAPability-QA. Based on the QA accuracy, we introduce a new metric, know but cannot tell ($K\bar{T}$), which evaluates the situation when a model knows the answer (i.e., can answer correctly when given it the related question), but cannot tell automatically in the caption without specific question as prompt. This evaluation is significant to the caption task of MLLMs, but is usually ignored by previous methods. Given the set of correct answers as $S_{qa}(\text{Correct})$, $K\bar{T}$ can be calculated as:
$K\bar{T}=\frac{\left|S_{qa}(\text{Correct})\cap(S(\text{Neg})\cup S(\text{Miss}))\right|}{|S_{qa}(\text{Correct})|}.$ (4)
@@
Table 4: The precision, recall, and F1-score of 7B open-source models on all dimensions. We keep their default settings for each model.
@@ Methods
F1-score LLaVA-OV-7B 78.1 34.3 66.0 62.7 88.5 37.1 74.3 83.9 8.7 47.6 37.3 15.5 41.3 52.0
Qwen2VL-7B 79.4 31.7 64.3 60.7 89.7 39.5 83.0 77.1 7.6 24.5 41.2 20.3 41.8 50.8
NVILA-8B 79.4 35.0 66.3 61.0 87.8 40.9 79.2 78.2 12.5 17.1 28.9 9.6 32.6 48.3
InternVL2.5-8B 74.9 33.3 64.9 63.5 88.6 42.5 81.5 81.9 16.3 31.2 36.2 19.7 41.8 52.0
VideoLLaMA3-7B 78.9 33.9 65.7 65.3 89.4 41.7 77.0 80.2 8.0 13.3 42.3 27.4 45.3 51.4
Qwen2.5VL-7B 80.6 31.1 68.5 63.1 90.4 47.5 87.1 82.0 16.4 31.9 42.0 26.6 44.9 54.8
Figure 7: The evaluation of repeating 5 times for Gemini-1.5-pro captions. We tag the fluctuation range beside the data point.
4 Experiments
4.1 Experimental Setups
For comprehensively evaluating the state-of-the-art (SOTA) models, we both choose several popular open-source and closed-source MLLMs. For closed-source models, we evaluate GPT-4o (0806) [27], Gemini-1.5-pro [30], and Gemini-2.0-flash [12]. For open-source models, we evaluate InternVL2.5 [9], LLaVA-OneVision [19], NVILA [24], VideoLLaMA3 [44], Qwen2VL [36] and Qwen2.5VL [32] with their different LLM sizes. We use the same image prompt and video prompt to infer all models, see Sec. B.2 for the inference prompts. We use GPT-4-Turbo-128k [1] to take scores for all generated captions to complete our evaluation. See Sec. A.1 for more implementation details.
4.2 Main Evaluation Results
F1-score of closed-source API and 72B models. We report the precision, recall, and F1-score of closed-source and 72B models in Tab. 3. Gemini-2.0-flash and GPT-4o achieve the highest precision (79.3% and 79.2%), which represents their captions are truthful and accurate. When it comes to the recall metric, Gemini-1.5-pro achieves the best with 3.9% higher than second place, i.e., GPT-4o, which means it is better at identifying more elements correctly. This leads the F1-score still ranked first by Gemini-1.5-pro (66.8%), and followed by GPT-4o and Gemini-2.0-flash. Gemini-1.5-pro has a huge advantage in object counting in both image (7.4% better than the second place) and video (16.4% better than the second place). GPT-4o excels at recognizing camera angle, as it is 9.6% higher than the second place. Qwen2.5-VL performs the best in the open-source models as it performs well on scene and camera movement. It is worth noting that these models behave differently in different dimensions. Object category, scene, OCR, and style seem simple for these powerful models as they all achieve well on the F1-score. When it comes to the dimensions of object number, object color, spatial relation, style, character identification, and events, all of them show relatively high precision but low recall, which means they can describe these elements well when they are confident about them, but might miss some instances and ignore the thoroughness. As for the action and two camera-related dimensions, all models perform unsatisfactorily on both precision and recall. This phenomenon inspires researchers to focus more on these aspects of the model’s capability.

### v3

@@ Methods
Question-answer pairs evaluation. As we annotate each descriptive element for each dimension rather than caption sentence, we can also convert our annotations to question-answer (QA) pair format to evaluate the MLLMs’ general ability out of the horizon of caption only. See Appendix A.5 for examples of our CAPability-QA. Based on the QA accuracy, we introduce a new metric, know but cannot tell ($K\bar{T}$), which evaluates the situation when a model knows the answer (i.e., can answer correctly when given it the related question), but cannot tell automatically in the caption without specific question as prompt. This evaluation is significant to the caption task of MLLMs, but is usually ignored by previous methods. Given the set of correct answers as $S_{qa}(\text{COR})$, $K\bar{T}$ can be calculated as:
$K\bar{T}=\frac{\left|S_{qa}(\text{COR})\cap[S(\text{INC})\cup S(\text{MIS})]\right|}{|S_{qa}(\text{COR})|}.$ (3)
4 Experiments
4.1 Experimental Setups
For comprehensively evaluating the state-of-the-art (SOTA) models, we both choose several popular open-source and closed-source MLLMs. For closed-source models, we evaluate GPT-4o (0806) gpt4o , Gemini-1.5-pro gemini1.5 , and Gemini-2.0-flash gemini2.0 . For open-source models, we evaluate InternVL2.5 internvl2.5 , LLaVA-OneVision llavaov , NVILA nvila , VideoLLaMA3 videollama3 , Qwen2VL qwen2vl and Qwen2.5VL qwen2.5vl with their different LLM sizes. We use the same image prompt and video prompt to infer all models, see Appendix B.2 for the inference prompts. We use GPT-4 Turbo (1106-preview) gpt4 to take scores for all generated captions to complete our evaluation. See Appendix B.1 for more implementation details.
4.2 Main Evaluation Results
Precision and hit of closed-source API and 72B models. We report the precision and hit of closed-source and 72B models in Tab. 3. Gemini-2.0-flash and GPT-4o achieve the highest precision (79.3% and 79.2%), which represents their captions are truthful and accurate. When it comes to the hit metric, all results drop significantly as it is harder to cover as many visual elements as possible. Gemini-1.5-pro achieves the best with 3.9% higher than second place, i.e., GPT-4o, which means it is better at identifying more elements correctly. Among 72B models, Qwen2.5VL achieves the best hit, with notable precision. When we focus on each dimension, it is worth noting that these models behave differently in different dimensions. Gemini-1.5-pro has a huge advantage in object counting in both image and video, especially in thoroughness (10% better than the second place for image and 16.3% better than the second place for video in hit metric). GPT-4o excels at recognizing camera angle, as it is 9.6% higher than the second place on both precision and hit. Qwen2.5-VL performs the best hit in the open-source models as it performs well on scene and camera movement. Object category, scene, OCR, and style seem simple for these powerful models, as they all achieve well on the F1-score. When it comes to the dimensions of object number, object color, spatial relation, style, character identification, and events, all of them show relatively high precision but low hit, which means they can describe these elements well when they are confident about them, but might miss some instances and ignore the thoroughness. As for the action and two camera-related dimensions, all models perform unsatisfactorily on both precision and hit. This phenomenon inspires researchers to focus more on these aspects of the model’s capability.
@@ 4.2 Main Evaluation Results
Table 4: The precision and hit of 7B open-source models on all dimensions. We keep their default settings for each model.
@@ Methods
Table 5: The $\uparrow$ (higher is better) of CAPability-QA and the $K\bar{T}\downarrow$ (lower is better) result.
Methods
Obj.
Cate.
Obj.
Num.
Obj.
Color
Spa.
Rel.
Scene
Cam.
Ang.
OCR Style
Cha.
Iden.
(D) Obj.
Num.
Act.
Cam.
Mov.
Event Avg.
QA Acc LLaVA-OV-72B 95.0 54.6 63.8 94.0 96.2 60.6 66.3 82.0 32.1 52.2 75.5 15.7 85.3 67.2
Qwen2VL-72B 94.7 56.1 68.6 90.7 94.0 65.0 82.4 86.6 31.3 48.9 73.0 34.1 72.7 69.1
InternVL2.5-78B 95.5 56.9 67.1 90.0 91.2 54.1 79.5 82.5 19.1 49.7 79.1 23.3 81.7 66.9
Qwen2.5VL-72B 92.7 58.2 67.4 84.4 88.7 63.9 87.4 87.3 33.4 41.4 75.8 39.5 85.8 69.7
GPT-4o (0806) 94.5 47.2 72.5 79.5 84.5 71.6 80.5 79.3 37.2 46.2 81.1 20.5 78.6 67.2
Gemini-1.5-pro 97.3 51.6 78.8 94.4 87.1 56.8 84.8 84.2 41.2 51.2 74.4 32.2 82.8 70.5
Gemini-2.0-flash 98.3 46.8 73.3 93.4 95.2 57.6 84.8 74.5 49.1 44.2 81.6 24.8 86.6 70.0
$K\bar{T}$ LLaVA-OV-72B 20.3 64.3 39.6 49.6 13.9 33.0 13.7 9.4 74.8 66.1 53.2 31.4 67.1 41.3
Qwen2VL-72B 16.7 62.1 37.0 46.0 12.6 35.4 10.2 10.4 84.7 78.3 47.6 53.1 60.4 42.6
InternVL2.5-78B 19.1 57.8 35.4 45.2 11.4 21.4 11.0 47.0 8.2 73.0 62.4 68.1 69.9 40.8
Qwen2.5VL-72B 15.3 60.7 33.5 37.2 9.1 24.3 5.9 8.5 47.4 66.2 49.3 38.2 61.1 35.1
GPT-4o (0806) 13.1 55.2 26.6 34.7 7.8 16.1 6.7 3.5 30.9 64.8 41.7 53.9 51.7 31.3
Gemini-1.5-pro 11.9 41.0 24.1 36.4 9.6 19.2 5.5 3.1 18.0 51.6 36.1 23.4 47.9 25.2
Gemini-2.0-flash 16.3 52.6 32.8 45.1 12.6 9.0 4.5 3.2 25.7 58.4 45.9 23.1 54.5 29.5
Precision and hit of 7B models. The precision and hit of 7B models are shown in Tab. 4. Among these 6 models, Qwen2.5VL-7B achieves the best precision (75.5%) and hit (48.1%), demonstrating its awesome ability. Averagely, the 7B models perform a bit worse than 72B models, verifying the scaling law. Among the dimensions, they follow a similar pattern as 72B and closed-source models. Researchers should focus on the thoroughness of object number, object color, spatial relation, style, character identification, and event, and try to improve the ability of the action and the two camera-related dimensions.
QA-based evaluation and the $\bm{K\bar{T}}$ metric. Due to the visual element annotation type, we can easily convert our annotations to QA format, evaluate the accuracy of closed-source APIs and 72B models, thus further calculating their $K\bar{T}$ metric, as shown in Tab. 5. We are surprised to find that the difference in their QA accuracy is not significant, which means they can have a similar level of understanding of the correct visual descriptions based on the specific questions. When it comes to the $K\bar{T}$ metric, these models vary a lot. Gemini-1.5-pro performs the best with the lowest $K\bar{T}$ (25.7%), but the $K\bar{T}$ of LLaVA-OneVision-72B, Qwen2VL-72B, and InternVL2.5-78B comes to more than 40%, which means they are more likely knowing the answer but cannot tell automatically. This phenomenon shows the performance gap between the strong instruction (QA) task and the weak instruction (caption) task, which may be ignored by previous work.
4.3 Metric Analysis
Consistency with human evaluation. To demonstrate the reliability of our annotations and evaluation pipeline, we build a human arena to manually compare the caption performance of closed-source models and 72B open-sourced models for each dimension. Specifically, we randomly select 5K pair-wise caption samples for each dimension and require human annotators to select a better one each time. The judging criteria include both the correctness and thoroughness for each dimension. We count the model ranking of each dimension, and compare it with the harmonic mean of precision and hit. The results are shown in Fig. 7. On the scene dimension, scores of all models are high and close in our evaluation, which may cause slight inconsistencies. But for most dimensions, our ranking shows high consistency with human results, leading to high reliability of final evaluation results.
Figure 7: The metric ranking consistency comparison with human arena evaluation.
Figure 8: Consistency comparison with other metrics, we calculate CIDER, BLEU_4, METEOR metrics, and conduct direct GPT-4-Turbo evaluation. Our evaluation shows the highest consistency.
Comparison with other metrics. We compare our metrics with other traditional metrics to demonstrate the credibility of our CAPability. We randomly select 200 images or videos for each dimension, and ask human annotators to annotate truthful and detailed captions as the ground truth. Then we calculate traditional metrics (i.e., CIDER vedantam2015cider , BLEU_4 papinesi2002bleu , and METEOR banerjee2005meteor ) between the human-annotated and model-generated captions. Besides, we also use GPT-4-Turbo gpt4 to directly score the captions from 0 to 5 based on the ground truth. We further randomly select 200 inference samples for each dimension and each model, and manually score them to get human evaluation results as a reference. We report the harmonic mean for CAPability and linearly scale all results to 0-1 for easier comparison. The results are shown in Fig. 8. As our result shows high consistency with human evaluation, all traditional metrics fail to align. The phenomenon shows that existing traditional metrics cannot reliably and reasonably evaluate the detailed visual captions. The direct GPT4-Turbo evaluation shows better consistency than traditional metrics, but still has gaps with CAPability, lacks reasonable interpretability, and is too vague for evaluation. Our CAPability can give scores from various aspects with clear judging criteria, thus obtaining an accurate and comprehensive evaluation.
5 Conclusion
In this work, we present CAPability, the first comprehensive visual caption benchmark through 6-view and 12-dimensional analysis. Unlike existing benchmarks that rely on oversimplified metrics or limited visual elements, CAPability introduces a correctness-thoroughness dual evaluation framework based on precision, hit, and $K\bar{T}$. Through this meticulous evaluation process, we uncover specific areas needing improvement across leading models, such as their gap between precision and hit, the challenges in aspects like camera angle detection, character identification, and action recognition. We also indicate the "know but cannot tell" phenomenon of MLLMs, which may be ignored by previous work. We believe that CAPability will play a pivotal role in advancing the field of visual captioning by encouraging the development of models that holistically understand and describe visual content. We open-source all our annotated data to facilitate future research.
Appendix
[…]
Appendix A More Details of CAPability
A.1 Details of Dimension Design
We argue that multi-dimensional evaluation is significant to visual caption evaluation and is more comprehensive than previous work. So how to choose proper dimensions? We refer to existing VQA benchmarks mmbench ; seedbench ; mvbench ; videomme and visual generation benchmarks geneval ; vbench ; t2vcompbench . VQA benchmarks usually design various types of questions to include multi-dimensional evaluation and analysis of MLLMs. For instance, MMBench mvbench defines 20 ability dimensions, including attribute recognition, attribute comparison, action recognition, spatial relationship, physical property, OCR, object localization, image style, image scene, identity reasoning, etc. MVBench mvbench covers 20 challenging video tasks including action, object, position, count, scene, pose, attribute, character, cognition, etc. Due to the flexible design of questions, VQA benchmarks can be naturally built with comprehensive dimensions. Different from the VQA task, the visual caption task does not require specific questions, but inspects the alignment of visual and textual information. Visual generation is the inverse task of visual captioning, as it requires models to generate specific visual content based on detailed textual descriptions. GenEval geneval designs 6 different tasks to evaluate text-to-image alignment, including single object, two object, counting, colors, position, and attribute binding. VBench vbench comprises 16 dimensions, including subject consistency, background consistency, object class, human action, color, spatial relationship, scene, style, etc. We follow their explored dimensions to design proper dimensions for visual captioning. Finally, we design 6 views, covering object, global, text, camera, temporal, and knowledge. The object-related view includes object category, object color, object number, and spatial relation, the global-related view includes scene and style, the text-related view evaluates the OCR capability of captions, the camera-related view covers the camera angle and movement, the temporal-related view contains action and event, and we also design a view to evaluate the knowledge of MLLMs, i.e., character identification.
We believe these dimensions contribute to a comprehensive visual caption benchmarking. However, it is undeniable that there may still be other dimensions that also contribute to caption evaluation. This phenomenon exists in all multi-dimensional benchmarks, but the purpose of our design is to find dimensions that are as comprehensive as possible and sufficient to differentiate model capabilities. As we design 12 dimensions, the evaluation is strong enough to evaluate models from various aspects. We also welcome the proposal of more constructive dimensions.
We explain each dimension in detail about what it represents here.
  • Object category. This dimension measures the ability of whether models can give a correct description of a specific object in the image. The object is randomly selected from the image.
  • Object color. Given a kind of randomly selected object in an image, this dimension measures the ability of whether models can correctly describe the color.
  • Spatial relation. Given two nearby objects in an image, this dimension measures the ability of whether models can correctly describe the spatial relationship of the two objects. We sample 500 images from our collected data, and sample another 500 images from CompreCap comprecap , with their spatial relationship descriptions.
  • Action. Given a video, this dimension measures the ability of whether models can recognize the action in the video. We use the video data of Dream-1K dream1k and re-annotate the action from their annotations.
  • Event. Given a video, this dimension measures the ability of whether models can tell a complete event in the video. We refer Dream-1K dream1k to design this dimension, and we extract the events from their annotations. Different from other dimensions with atom-level elements, the event is usually composed of subjects and actions, which measures the temporal summarization ability of the model.
A.2 Explanation for One Represents All Strategy
For each dimension, we only annotate one element, though there may be more than one element existing for some dimensions. Therefore, our annotations do not cover the whole visual content. But for those who try to cover the whole visual content, it is actually pretty hard to achieve this, as we mentioned in Sec. 3.1, the objects can be divided into almost infinite granularity. We focus on keeping the randomness of element selection, thus covering the whole visual content in a statistical sense, based on the law of large numbers. Therefore, we can get the ability to evaluate the thoroughness of the generated captions by calculating the hit.
Figure A1: Two examples (object number and camera angle) of data distribution before data balance (pre-annotated) and after data-balanced selection (final human-annotated).
A.3 Details of Data Balance and Final Distribution
The purpose of data balance is to suppress the long-tail distribution, thus ensuring there are a certain number of samples of different difficulties in the benchmark. Fig. A1 shows two examples of the comparison of pre-annotated data distribution and final human-annotated data distribution. For object number and camera angle dimension, we first randomly sample nearly 75K samples and conduct model pre-annotation. The 75K samples consist of approximately 1/3 SA-1B sam , 1/3 COYO-100M coyo700m , and 1/3 our crawled from multiple public datasets and websites. For object number dimension, we select out all samples with counting from 2 to 16 same objects within an image. As shown in Fig. A1 (left upper), the counting follows the long-tail distribution, there are fewer images with more objects within an image. Therefore, we conduct data-balanced sampling. Specifically, we separately sample images for different counts, thus forcing the number of each counts more balanced after the human correcting and filtering process. For camera angle dimension, the dutch angle data is rare, therefore we keep all dutch angle data, and sample the same number of other three categories. After the human correcting and filtering, the number of these categories varies slightly. The situation in other dimensions is also similar.
As we consider the counts, categories, etc. for each dimension to conduct the data balance and do not consider the data source (i.e., from SA-1B, COYO-700M or crawled by ourselves) during this process, the final data source distribution for dimensions of object category, object number, object color, scene, camera angle, OCR varies. For spatial relation, we directly choose 1/2 CompreCap comprecap and 1/2 SA-1B images as SA-1B is more likely to contain high-resolution images with complex object relationship scene. For style, we choose all realistic images from SA-1B, and crawl animated, special effect, and old-fashioned images by ourselves, all art-related images are from Wikipaintings wikipaintings . For character identification, we use all images from public dataset, i.e., Wukong-100M wukong100m rather than crawling to ensure the proper copyright. For dynamic object number, we directly use data from VSI-Bench vsibench . For action and event, we directly use videos from Dream-1K dream1k . We crawl all videos for camera movement dimension by ourselves as there are few camera movement data in existing datasets. We thank all public datasets and benchmarks, their excellent images, videos, and annotations provides much convenience for building our CAPability.
Figure A2: Bad case examples of pre-annotation or metadata, human annotators filter or correct the wrong pre-annotation or metadata one by one.
A.4 Discussion about the designed metrics
We define two metrics, precision and hit, to evaluate both the correctness and thoroughness. The hit is similar to the meaning of recall, which measures how many visual elements in the image/video can be described correctly. However, the recall is usually related to the TP/FP/FN/TN system allen1955machine , which is inconsistent with our evaluation situations. To avoid ambiguity and misunderstanding, we name our metric as hit rather than recall.
We give the analysis from the TP/FP/FN/TN perspective here. The inconsistency comes from the conflict of the situation definition. In our pipeline, there are three situations when evaluating: 1) MIS, 2) COR, 3) INC, but the TP/FP/FN/TN system does not define the situation of MIS. In the TP/FP/FN/TN system, the precision and recall are defined as:
Precision $=\frac{TP}{TP+FP},$ (A1)
Recall $=\frac{TP}{TP+FN}.$ (A2)
The existence of MIS leads to the ambiguity of the definition of $FN$ and $FP$. If there is no existing MIS, we can only calculate the accuracy without considering precision or recall. Now we try to analyze with MIS. Based on the definition of $TP$, we can map our COR to $TP$ with no doubt, as it correctly predicts the answer. If we want to calculate the precision, we can map our INC to $FP$, as it wrongly predicts the answer. Therefore, the precision can be calculated by Eq. 1 or Eq. A1, which is consistent. When we consider recall, $TP+FN$ should be the number of all ground truths, as there are no negative samples ($TN$) in our annotation, which means $TP+FN=S(\text{COR})\cup S(\text{INC})\cup S(\text{MIS})$. This leads to S(INC) being included in both $FP$ and $FN$. As the TP/FP/FN/TN system does not define the MIS, the TP/FP/FN/TN-based precision and recall cause contradiction. Therefore, we name our metric of $S(\text{COR})/(S(\text{COR})\cup S(\text{INC})\cup S(\text{MIS}))$ as hit rather than recall to avoid ambiguity as it does not fit the TP/FP/FN/TN-based definition.
In addition to the TP/FP/FN/TN system, there are also other ways to define precision and recall. MUC-7 chinchor1998appendix defines the precision and recall with the situation of MIS. Apart from COR, INC, MIS, which own the same meaning as ours, MUC-7 also defines SPU, which represents the number of spurious, and equals 0 in our situation. MUC-7 defines the precision and recall as follows:
Precision $=\frac{S(\text{COR})}{S(\text{COR})+S(\text{INC})+S(\text{SPU})},$ (A3)
Recall $=\frac{S(\text{COR})}{S(\text{COR})+S(\text{INC})+S(\text{MIS})}.$ (A4)
Based on this kind of definition, our hit equals the "recall".
However, as the TP/FP/FN/TN system is too famous and standard to define the precision and recall, we finally decide to use the "hit" rather than "recall" to avoid misunderstanding.
A.5 Benchmark Examples
Examples of human annotation process. We show some visual cases of the human annotation process in Fig. A2. All of examples are with wrong pre-annotation or metadata. Human annotators check the pre-annotation/metadata one by one and filter/correct the mistakes for each dimension.
Examples of converted QA pairs. As we directly annotate the visual elements in the image or video rather than the caption sentence, we can easily convert our annotation into the format of question-answer (QA) pairs, and we name it as CAPability-QA. We use CAPability-QA to evaluate the QA accuracy and the know but cannot tell ($K\bar{T}$) metric. In Fig. A4, we also show the same visual cases as Fig. A3 for each dimension with converted QA format. Most of the dimensions are converted to the format of a multiple-choice QA task with several options, and the object color, OCR, and character identification dimensions are designed as open-ended QA tasks.
Appendix B More Experimental Analysis
B.1 Implementation Details
We use 4x80G GPUs to run all open-sourced model inference. We use transformers to deploy LLaVA-OneVision, InternVL2.5, VideoLLaMA3, and NVILA, use vLLM to deploy Qwen2VL and Qwen2.5VL, as their official repositories suggested. For all our evaluated model, we follow their official configurations to run the inference. We set the temperature of all open-source models to 0, while keeping the default for closed-source APIs. All maximum output token length is set to 8192. We list the configurations as follows.
NVILA. We use the official image and video process function in VILA repository, and uniformly sample 8 frames for videos, as suggested in the official config.
Gemini-1.5-pro and Gemini-2.0-flash. As Gemini API supports video, we directly send the original image and video to the API server. For very few videos with too large file size, we downsample the fps to 3, and send the downsampled video to the API server for connection stability.
B.2 Prompts of Inference and Evaluation
Evaluation prompt. As we divide the evaluation of dimensions into two types: 1) dimensions with specific categories (i.e., style, camera angle, and camera movement), 2) dimensions with open-ended descriptions. Therefore, we design two kinds of templates for evaluating, and fine-tune them within each dimension. In Fig. A6, we take the object number dimension and camera movement dimension as examples, to show our prompts for evaluation. For dimensions with specific categories, we ask GPT-4-Turbo to extract the key information and classify the caption into our pre-defined categories or the ’N/A’ class. The correct classification is considered positive, the wrong one as negative, and the ’N/A’ result is considered a miss. For dimensions with open-ended descriptions, we ask GPT-4-Turbo to directly compare the annotations and the caption, and give out the result of positive, negative, or miss with reasons.
B.3 More Experimental Results